What is magnetic field due to toroid.


1
Expert's answer
2021-01-06T13:52:52-0500

"\int Bdl=\mu_0I\to B\cdot2\pi\cdot\frac{r+R}{2}=\mu_0NI\to"


"B=\frac{\mu_0NI}{\pi(r+R)}" . Answer